OpenStudy (anonymous):

The equation of the tangent line to f(x) = sqrt x at x = 49 is y = . Using this, we find our approximation for sqrt 49.3 is

OpenStudy (perl):

L(x) ~ f(x) as x->a , where L(x) = f (a) + f ' (a) * ( x - a) choose a = 49 L(x) = f(49) + f ' (49 ) ( x - 49 ) L(x) = sqrt(49) + 1/(2*sqrt(49)) * ( x - 49) L(x) = 7 + 1/(2*7) ( x - 49 )

OpenStudy (anonymous):

f(49) is 7?

OpenStudy (perl):

yes, because f(x) = sqrt(x)

OpenStudy (anonymous):

then what do you do with the x at the end?

OpenStudy (perl):

you can distribute it

OpenStudy (anonymous):

so the answer has an x in it?

OpenStudy (perl):

yes, since it is a tangent line

OpenStudy (anonymous):

so then whats y

OpenStudy (anonymous):

@perl